# NeuCo-Bench: A Novel Benchmark Framework for Neural Embeddings in Earth Observation

### v1

@@ 7 Conclusion
We presented NeuCo-Bench, the first of its kind, task-driven benchmarking framework for compression that evaluates neural embeddings by downstream task performance, rather than pixel fidelity. The framework introduces a novel rank-then-aggregate scoring method which dynamically determines the task complexity based on score statistics. We demonstrated the NeuCo-Bench framework by setting up the CVPR EARTHVISION 2025 data challenge on multi-modal, multi-temporal, and open-source EO data. We introduced a novel set of real-world downstream tasks which remained undisclosed at the time of the data challenge, and have been publicly released after the conclusion of the competition. NeuCo-Bench encourages the development of methods that generate semantically-rich, general-purpose embeddings.
For our setup, experiments demonstrated that multi-modal foundation models yields strong overall performance—particularly on semantic land-cover tasks. Post-encoding fusion of seasonal views resulted in notable gains for temporally sensitive tasks such as cloud cover prediction. We also observed that smaller and, in some cases, larger embedding sizes degrade performance. This observation highlights compact embeddings as a practical choice for image-level tasks when high-quality annotations and compute resource become scarce.
NeuCo-Bench is open source and ready for extension—either by novel evaluation methods or additional downstream tasks without any conceptual restriction to Earth observation. Currently, the NeuCo-Bench framework is limited to image-level tasks, but future work aims to extend NeuCo-Bench’s functionalities to include pixel-wise outputs, options beyond linear probing, and an assessments of bit-rate efficiency.
@@ Acknowledgment
Moreover, we are grateful for the active participation in our CVPR EARTHVISION 2025 workshop data challenge with valuable Q&A in https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement/issues?q=is%3Aissue. In particular, conversations with Burak Ekim (UniBW) and Isaac Corley (Wherobots) sparked lively discussions, and our meetings with the winning teams KTH and Friends (KTH Royal Institute of Technology, Sweden) and 404 Embedding Not Found (Microsoft Research, USA).
@@ A.1 Challenge Evaluation Method and Configuration
$Q_{t}^{(p)}=100\epsilon\frac{\left\langle s_{t,k}\right\rangle_{k}}{\text{std}_{k}\left(s_{t,k}\right)+\epsilon}\equiv Q=100\epsilon\frac{\bar{s}}{\Delta s+\epsilon}\quad\text{with}\quad\epsilon=0.02\quad.$ (3)
@@
13 [ s_sub for s_sub in q.values() if sign*s_sub < sign*s ]
@@
$s^{(p)}=\sum_{t=1}^{T}w_{t}R_{t}^{(p)}\quad\text{where}\quad w_{t}=\text{std}_{p}Q_{t}^{(p)}\Big/\sum_{t=1}^{T}\text{std}_{p}Q_{t}^{(p)}$ (4)
@@
$\delta_{0}=\sqrt{\left\langle Q_{0}^{(p)2}\right\rangle_{p}}>0\quad\text{setting}\quad\bar{Q}_{0}=0\quad\text{and defining}\quad R_{0}^{(p)}=0~.$ (8)
@@ A.2 Platform and Infrastructure
running on top of the Juelich Supercomputing Center’s (JSC) OpenStack [^5] https://www.openstack.org cloud environment. The communication with the Eval.AI API for fetching submission data and writing results back to the Eval.AI leaderboard was based on the Eval.AI GitHub remote challenge evaluation template [^6] https://github.com/Cloud-CV/EvalAI-Starters/commit/8338085c6335487332f5b57cf7182201b8499aad utilizing the requests Python library. [^7] https://docs.python-requests.org Figure˜6 illustrates the entire setup: (a) the Eval.AI web interface [^8] https://eval.ai/web/challenges/challenge-page/2465 and a supplementary repository [^9] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement on one end, and (b) the evaluation procedure which runs on the VM at JSC, on the other end.
As evaluation method (Evaluate.py), NeuCo-Bench [^10] https://github.com/embed2scale/NeuCo-Bench/tree/040ca567da4d231ced78a16448d2039a0e871276 was incorporated into the Eval.AI remote challenge evaluation template running on the VM, cf. Local Repository – Evaluation Codebase in Fig.˜6. Updates to the Eval.AI challenge web interface got triggered by GitHub Actions. [^11] https://docs.github.com/en/actions In addition, the Supplementary Repository [^12] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement serves two purposes:
@@ A.3 Competition Analysis
Figure 7: The evolution of participant rankings in the challenge test phase. Lines correspond to participating teams and dots to submissions by the corresponding team. Team 10 is the simple mean baseline described in Section˜B.3.1, and Team 11 is a randomized baseline with randomly sampled, normally distributed embeddings.
@@
In general, the ranking and the mean $Q$ value are close to identical. However, the team achieving third place upended the order of first and second place, with the effect that the runner-up team achieved a slightly higher mean $Q$-score than the winners. This effect is driven by a change in task weights caused by the third-place team’s performance. We note the dynamics around Submissions 14 and 15 as illustrated in Fig.˜7 where the ranking dynamics given a sequence of Submissions (dots) is documented: Team 1 through 9 are competing numerically indexed by their final position in the challenge ranking. Team 10 represents the simple mean baseline case described in Section˜5.3 with additional details in Section˜B.3.1, and Team 11 is a randomized baseline submitting randomly sampled, normally distributed embeddings.
@@
Figure 8: Visualisation of downstream-task labels (top row) and corresponding Sentinel-2 images (bottom row). The biomass labels have been sourced from Karaman et al. (2025). Although Sentinel-1 is included in every data cube, it was excluded from this visualisation. The cloud and heat island labels are based on aggregated images.
Since the competition concluded in early April 2025, the one winner and the fourth-place team have open-sourced their solutions: the team achieving highest mean $Q$ is available, [^14] https://github.com/KerekesDavid/embed2scale-solution and the fourth-place team, who avoided (pre-)training completely by utilizing MOSAIKS (Rolf et al., 2021) is available, [^15] https://github.com/isaaccorley/temporal-mosaiks too.
@@ Appendix B An Extendable Framework
Based on our insights from the CVPR EARTHVISION 2025 data challenge, we took our approach to the next level with the intention to build a community around benchmarking neural compression. Table˜2 provides a high-level comparison on how NeuCo-Bench fits into existing, popular geospatial benchmarking frameworks. In summary, NeuCo-Bench fills the following gaps:
@@
It is worth iterating that our dataset SSL4EO-S12-downstream dataset builds on the SSL4EO-S12 data structures part of GEO-Bench—thus, serving as a potential interface regarding synergies. We are currently forming the Earth2Vec community around NeuCo-Bench where more than 30 organizations from academia, the corporate world, and governments have joined.
@@ B.1 Benchmark Tasks
Figure˜8 illustrates examples of Sentinel-2 inputs alongside the corresponding labels. Table˜3 lists a detailed overview of the derived 11 tasks, whereby 9 of these were used in the CVPR EARTHVISION 2025 data challenge and are highlighted with green check marks. Each task is linked to a HuggingFace identifier, following the definition in SSL (2025) and listed in Table˜3. Clouds and Nodata were not included in the CVPR competition but are provided with the release of NeuCo-Bench. The task Random provides randomly generated labels and associated data cubes from the Cloud task, which introduced an additional quality assessment.
The processing pipeline for all presented datasets utilized GEE GEE to download data cubes, applying a maximum cloud coverage filter of 10%, as provided by the GEE property CLOUD_COVER, except for the clouds task where no restrictions on cloud cover were enforced. Each data cube was aligned to the center of the corresponding label and processed to a size of 264 x 264 pixels. All sample locations with less than 4 images were discarded, following the requirements in Blumenstiel et al. (2025). Whenever possible, only locations that cover all four seasons for Sentinel-1 and Sentinel-2 were chosen. In case of missing latitude and longitude, locations were randomly selected from shapefiles representing regions such as mainland Europe or areas within the US Corn Belt. Comparison between the presented dataset and the data provided in Blumenstiel et al. (2025)show one major difference. In January 2022, ESA introduced a new baseline for Sentinel-2 data, effectively shifting all pixel intensities by 1000 units upward. The dataset presented in this work follows the format of GEE, i.e. removing this translation such that the minimum value for Sentinel-2 pixels are 0 both before and after the change by ESA. Blumenstiel et al. (2025) adheres to the ESA standard and enforces a lower bound for Sentinel-2 pixels of 1000, before and after the change. To allow for seamless integration between the two datasets, the dataloaders provided in NeuCo-Bench includes a setting that toggles a shift by 1000, aligning the distributions of the two datasets.
@@
Biomass biomass_mean__regr Regression tasks: Biomass density (Mg/ha) mean and standard deviation estimated for pixel-level labels derived from GEDI.
@@
Crops crops__regr Regression task: Combined fraction of Soybean and Corn in the label image.
Landcover landcover_agriculture__regr Regression task: Percentage of agriculture pixels in the Corine Land cover image.
@@
Clouds clouds_reg__regr Regression task: Average cloud cover fraction across four seasons in one year.
Heatisland heatisland_mean__regr Regression tasks: Summer surface temperature mean and standard deviation in Kelvin.
@@
Public vs. Secret Downstream Tasks. We released the hidden NeuCo-Bench downstream tasks (cf. Table˜3 with green check mark) after the conclusion of the 2025 CVPR EARTHVISION workshop to make publicly accessible the standalone SSL4EO-S12-downstream dataset for reasons of transparency, and to be used and contributed to by the neural compression community. As common with public benchmarks designed for standalone usage, we assume that NeuCo-Bench users would not jeopardize the developments of their own compressor $E$ by willingfully exploiting knowledge of the downstream tasks they test on. Removing and adding (mix-and-match) downstream tasks in NeuCo-Bench for a new competition avoids overfitting of a state-of-the-art compressor $E$. The process is as straightforward as uploading such data $x$ to HuggingFace, i.e.,
@@ B.2 Standalone Python Implementation
In order for a clean separation of code from the open-source platform Eval.AI, we developed a minimal viable standalone Python code base [^16] https://github.com/embed2scale/NeuCo-Bench/tree/040ca567da4d231ced78a16448d2039a0e871276 to serve as plug-and-play for any larger ecosystem integrating of the core NeuCo-Bench framework. In fact, as Fig.˜6 demonstrates, the scoring for the Eval.AI leaderboard is entirely taken care of by NeuCo-Bench. Correspondingly, our framework commits an additional, customized leaderboard that globally depends on all submissions to a dedicated GitHub repository. [^17] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement/tree/272c911e8fb527d265de9be18fc81e16208ca0b6?tab=readme-ov-file#leaderboard
@@
6 --method_name ’your-method-name’ \
@@
where your-method-name and phase-name are free strings to define an output (sub-)directory
/path/to/results_directory/phase-name/your-method-name_YYYYMMDD_HHmmSS
@@
with YYYYMMDD a date of year YYYY and zero-padded numerical month MM and day DD. HHmmSS indicates a time of the day in hours HH, minutes mm, and seconds SS, accordingly. A YAML file /path/to/config.yaml specifies details of the evaluation such as:
1embedding_dim: 1024 # number of embedding dimensions
2batch_size: 64 # batch size for (linear) probing
3epochs: 20 # number of epochs to optimize the (linear) probe for
4learning_rate: 0.001 # learning rate to optimize with
5k_folds: 40 # number of cross-validations to generate statistics over
6standardize_embeddings: true # standardize embeddings by their global mean and std
7normalize_labels: true # normalize in range [0,1]
8task_filter: false # all in /path/to/annotation_directory/ per default
9 # example: ["biomass_mean", "biomass_std"]
10# etc.
B.3 General Evaluation Setup
We expand on the methods introduced in Section˜5.3, applying off-the-shelf embedding methods to the downstream tasks defined in Section˜B.1. Unlike the CVPR challenge, where data were partitioned into development and evaluation phases, results here are computed on the entire downstream dataset. Unless otherwise stated, evaluation follows the challenge protocol and used hyperparameters: $E=20$ epochs, $k=50$ training and test set splits, and a learning rate of $10^{-4}$. We report raw $R^{2}$ scores, clipping negatives to $[0,1]$ for figures.
For the results in Section B.3.1, we use embeddings of size $N=1024$, consistent with the challenge setup. In Section B.3.3, we relax this requirement to investigate the effect of varying embedding dimensionality and explore decoder choices beyond the linear probes proposed as default.

### v2

@@ 7 Conclusion
We present NeuCo-Bench, a task-driven benchmarking framework for evaluating compact embeddings by downstream task performance. The framework introduces a novel rank-then-aggregate scoring method which dynamically determines the task complexity based on score statistics. We demonstrated the NeuCo-Bench framework by running the 2025 CVPR EarthVision data challenge on multi-modal, multi-temporal, and open-source EO data and introducing a set of real-world downstream tasks which remained undisclosed at the time of the data challenge, and have been publicly released after the conclusion of the competition. NeuCo-Bench encourages the development of methods that generate semantically-rich, general-purpose embeddings.
We demonstrate that modern EO FMs, such as the multi-modal model TerraMind yield strong overall performance—particularly on semantic land-cover tasks. Post-encoding fusion of seasonal views results in notable gains for temporally sensitive tasks such as cloud cover prediction. We also observe that smaller and, in some cases, larger embedding sizes may degrade performance. This observation highlights compact embeddings as a practical choice for image-level tasks when high-quality annotations and compute resource become scarce.
NeuCo-Bench is open source and ready for extension—either by novel evaluation methods or additional downstream tasks without any conceptual restriction to Earth observation. Currently, NeuCo-Bench is limited to image-level tasks, but future work aims to extend the functionality to include pixel-wise outputs, options beyond linear probing, and an assessments of bit-rate efficiency.
@@ Acknowledgment
Moreover, we are grateful for the active participation in our 2025 CVPR EarthVision workshop data challenge with valuable Q&A in https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement/issues?q=is%3Aissue. In particular, conversations with Burak Ekim (UniBW) and Isaac Corley (Wherobots) sparked lively discussions, and our meetings with the winning teams KTH and Friends (KTH Royal Institute of Technology, Sweden) and 404 Embedding Not Found (Microsoft Research, USA).
@@ A.1 Challenge Evaluation Method and Configuration
$Q_{t}^{(p)}=100\epsilon\frac{\left\langle s_{t,k}\right\rangle_{k}}{\text{std}_{k}\left(s_{t,k}\right)+\epsilon}$ $\equiv Q=100\epsilon\frac{\bar{s}}{\Delta s+\epsilon}$ (3)
$\text{with}\quad\epsilon$ $=0.02.$
@@
13 [ s_sub for s_sub in q.values()
14 if sign*s_sub < sign*s ]
@@
$s^{(p)}$ $=\sum_{t=1}^{T}w_{t}R_{t}^{(p)}$ (4)
$\text{where}\quad w_{t}$ $=\frac{\text{std}_{p}Q_{t}^{(p)}}{\sum_{t=1}^{T}\text{std}_{p}Q_{t}^{(p)}}$
@@
$\delta_{0}$ $=\sqrt{\left\langle Q_{0}^{(p)2}\right\rangle_{p}}>0$ (8)
$\text{setting}\quad\bar{Q}_{0}$ $=0\quad\text{and defining}\quad R_{0}^{(p)}=0~.$
@@ A.2 Platform and Infrastructure
running on top of the Juelich Supercomputing Center’s (JSC) OpenStack [^5] https://www.openstack.org cloud environment. The communication with the Eval.AI API for fetching submission data and writing results back to the Eval.AI leaderboard was based on the Eval.AI GitHub remote challenge evaluation template [^6] https://github.com/Cloud-CV/EvalAI-Starters/commit/8338085c6335487332f5b57cf7182201b8499aad utilizing the requests Python library. [^7] https://docs.python-requests.org Figure˜7 illustrates the entire setup: (a) the Eval.AI web interface [^8] https://eval.ai/web/challenges/challenge-page/2465 and a supplementary repository [^9] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement on one end, and (b) the evaluation procedure which runs on the VM at JSC, on the other end.
As evaluation method (Evaluate.py), NeuCo-Bench [^10] https://github.com/embed2scale/NeuCo-Bench/tree/040ca567da4d231ced78a16448d2039a0e871276 was incorporated into the Eval.AI remote challenge evaluation template running on the VM, cf. Local Repository – Evaluation Codebase in Fig.˜7. Updates to the Eval.AI challenge web interface got triggered by GitHub Actions. [^11] https://docs.github.com/en/actions In addition, the Supplementary Repository [^12] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement serves two purposes:
@@ A.3 Competition Analysis
In general, the ranking and the mean $Q$ value are close to identical. However, the team achieving third place upended the order of first and second place, with the effect that the runner-up team achieved a slightly higher mean $Q$-score than the winners. This effect is driven by a change in task weights caused by the third-place team’s performance. We note the dynamics around Submissions 14 and 15 as illustrated in Fig.˜8 where the ranking dynamics given a sequence of Submissions (dots) is documented: Team 1 through 9 are competing numerically indexed by their final position in the challenge ranking. Team 10 represents the simple mean baseline case described in the main paper with additional details in Section˜C.2, and Team 11 is a randomized baseline submitting randomly sampled, normally distributed embeddings.
Figure 8: The evolution of participant rankings in the challenge test phase. Lines correspond to participating teams and dots to submissions by the corresponding team. Team 10 is the simple mean baseline described in Section˜C.2, and Team 11 is a randomized baseline with randomly sampled, normally distributed embeddings.
Task (# samples) $N=1024$ $N=512$ $N=2048$
Biomass (2415) 4.23 4.24 7.98 16.01 16.54 3.87 7.54 4.25 8.32
Crops (3355) 5.60 5.56 11.26 22.15 22.00 5.09 10.35 5.94 12.34
Clouds (1140) 2.04 2.03 3.94 7.76 7.59 1.91 3.69 2.13 4.12
Landcover Forest (4691) 7.86 7.81 15.48 31.06 30.47 7.08 14.37 8.06 16.64
Heatisland (1659) 2.79 2.80 5.53 10.84 10.83 2.66 5.25 2.89 5.85
No-data (13260) 22.45 22.45 44.30 88.68 88.18 20.46 42.03 23.18 46.60
@@
Since the competition concluded in early April 2025, the one winner and the fourth-place team have open-sourced their solutions: the team achieving highest mean $Q$ is available, [^14] https://github.com/KerekesDavid/embed2scale-solution and the fourth-place team, who avoided (pre-)training completely by utilizing MOSAIKS [54] is available, [^15] https://github.com/isaaccorley/temporal-mosaiks too.
A.4 Compute Resources & Tuning Parameters
  • Embedding dimension ($N$, embedding_dim)
  • Number of epochs per CV fold ($E$, epochs)
  • Number of CV folds ($k$, k_folds)
  • Choice of tasks included (task_filter)
Empirical runtime measurements confirm an approximately linear scaling w.r.t. both, the number of epochs $E$ and the number of cross-validation folds $k$. A similar scaling behavior was numerically verified for the dataset size (# samples) for fixed downstream task. Runtimes are further influenced–—though to lesser extent—–by the embedding dimensionality $N$. For example, increasing the embedding size from 512 to 1024 dimensions results in a runtime increase of approximately 5% to 10% across tasks. For $N=1024$ to $N=2048$ dimensions implies an additional increase of about 5% to 15%—depending on the task dataset size. Such a sub-linear scaling may be attributed to computation overheads and system-level inefficiencies. Those may reduce the relative computational costs when increasing the embedding dimensionality $N$. Table˜3 lists a collection of recorded execution times (in seconds) for various parameter configurations per downstream task. All runtimes were measured on a single commodity ARM64 CPU with 16 cores (4.06 GHz) and 64 GB of RAM.
@@ Appendix B An Extendable Framework
Based on our insights from the CVPR EARTHVISION 2025 data challenge, we took our approach to the next level with the intention to build a community around benchmarking compact neural embeddings. Table˜4 provides a high-level comparison on how NeuCo-Bench fits into existing, popular geospatial benchmarking frameworks. In summary, NeuCo-Bench fills the following gaps:
@@ B.1 Benchmark Tasks
Figure 9: Visualisation of downstream-task labels (top row) and corresponding Sentinel-2 images (bottom row). Although Sentinel-1 is included in every data cube, it was excluded from this visualisation. The cloud and heat island labels are based on aggregated images.
Figure˜9 illustrates examples of Sentinel-2 inputs alongside the corresponding labels. Table˜5 lists a detailed overview of the derived 11 tasks, whereby 9 of these were used in the CVPR EARTHVISION 2025 data challenge and are highlighted with green check marks. Each task is linked to a HuggingFace identifier, following the definition in [2] and listed in Table˜5. Clouds and Nodata were not included in the CVPR competition but are provided with the release of NeuCo-Bench. The task Random provides randomly generated labels and associated data cubes from the Cloud task, which introduced an additional quality assessment.
The processing pipeline for all presented datasets utilized GEE [1] to download data cubes, applying a maximum cloud coverage filter of 10%, as provided by the GEE property CLOUD_COVER, except for the clouds task where no restrictions on cloud cover were enforced. Each data cube was aligned to the center of the corresponding label and processed to a size of 264 x 264 pixels. All sample locations with less than 4 images were discarded, following the requirements in [8]. Whenever possible, only locations that cover all four seasons for Sentinel-1 and Sentinel-2 were chosen. In case of missing latitude and longitude, locations were randomly selected from shapefiles representing regions such as mainland Europe or areas within the US Corn Belt. Comparison between the presented dataset and the data provided in [8]show one major difference. In January 2022, ESA introduced a new baseline for Sentinel-2 data, effectively shifting all pixel intensities by 1000 units upward. The dataset presented in this work follows the format of GEE, i.e. removing this translation such that the minimum value for Sentinel-2 pixels are 0 both before and after the change by ESA. Blumenstiel et al. [8] adheres to the ESA standard and enforces a lower bound for Sentinel-2 pixels of 1000, before and after the change. To allow for seamless integration between the two datasets, the dataloaders provided in NeuCo-Bench includes a setting that toggles a shift by 1000, aligning the distributions of the two datasets.
@@
Biomass biomass_mean__regr Regression tasks: Biomass density (Mg/ha) mean and standard deviation estimated for pixel-level labels derived from GEDI [20] and sampled based on [55].
@@
Crops crops__regr Regression task: Combined fraction of Soybean and Corn in the label image [9].
Landcover landcover_agriculture__regr Regression task: Percentage of agriculture pixels in the Corine Land cover image [23].
@@
Clouds clouds_reg__regr Regression task: Average cloud cover fraction across four seasons in one year [4].
Heatisland heatisland_mean__regr Regression tasks: Summer surface temperature mean and standard deviation in Kelvin based on LandSat-8 [48].
@@
Public vs. Secret Downstream Tasks. We released the hidden NeuCo-Bench downstream tasks (cf. Table˜5 with green check mark) after the conclusion of the 2025 CVPR EARTHVISION workshop to make publicly accessible the standalone SSL4EO-S12-downstream dataset for reasons of transparency, and to be used and contributed to by the neural compression community. As common with public benchmarks designed for standalone usage, we assume that NeuCo-Bench users would not jeopardize the developments of their own compressor $E$ by willingfully exploiting knowledge of the downstream tasks they test on. Removing and adding (mix-and-match) downstream tasks in NeuCo-Bench for a new competition avoids overfitting of a state-of-the-art compressor $E$. The process is as straightforward as uploading such data $x$ to HuggingFace, i.e.,
@@ B.2 Standalone Python Implementation
In order for a clean separation of code from the open-source platform Eval.AI, we developed a minimal viable standalone Python code base [^16] https://github.com/embed2scale/NeuCo-Bench/tree/040ca567da4d231ced78a16448d2039a0e871276 to serve as plug-and-play for any larger ecosystem integrating of the core NeuCo-Bench framework. In fact, as Fig.˜7 demonstrates, the scoring for the Eval.AI leaderboard is entirely taken care of by NeuCo-Bench. Correspondingly, our framework commits an additional, customized leaderboard that globally depends on all submissions to a dedicated GitHub repository. [^17] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement/tree/272c911e8fb527d265de9be18fc81e16208ca0b6?tab=readme-ov-file#leaderboard
@@
6 --method_name ’method-name’ \
@@
where method-name and phase-name are free strings to define an output (sub-)directory phase-name/method-name_YYYYMMDD_HHmmSS under /path/to/results_directory/, with YYYYMMDD a date of year YYYY and zero-padded numerical month MM and day DD. HHmmSS indicates a time of the day in hours HH, minutes mm, and seconds SS, accordingly. A YAML file /path/to/config.yaml specifies details of the evaluation such as:
1# number of embedding dimensions
2embedding_dim: 1024
3# batch size for (linear) probing
4batch_size: 64
5# number of epochs to optimize the (linear) probe for
6epochs: 20
7# learning rate to optimize with
8learning_rate: 0.001
9# number of splits to generate statistics over
10k_folds: 40
11# standardize embeddings by their global mean and std
12standardize_embeddings: true
13# normalize in range [0,1]
14normalize_labels: true
15# Filter to include only specific tasks.
16# all in /path/to/annotation_directory/ per default
17# example: ["biomass_mean", "biomass_std"]
18task_filter: false
19
20# etc.
B.3 Licenses for Data & Software
Table 6: List of licenses related to datasets currently included in our benchmark. All of these, except Clouds, are available in GEE. the main paper lists years of target labels, ranging from 2018 through 2024.
Dataset Origin of Data License
Sentinel-1 & -2 ESA / Copernicus CC BY-SA 3.0 IGO
Landsat-8 USGS [48] Public Domain
CDL USDA NASS Cropland Data Layers [9] Public Domain
Copernicus Land Monitoring Service [23] and Free Access
CloudSen12+ CloudSEN12 project [4] CC0 1.0
GEDI NASA [20] Public Domain
NeuCo-Bench builds on open-source software and is released under the Apache 2.0 license publicly available at https://github.com/embed2scale/NeuCo-Bench. All package dependencies are listed in the requirements.txt [^25] https://github.com/embed2scale/NeuCo-Bench/blob/main/requirements.txt file, and those are licensed under widely-accepted open-source terms [^26] https://opensource.org/licenses, including BSD, MIT, PSFL, The Unlicense272727code-equivalent to CC0 data licenses, MPL-2.0282828weak copy-left that allows for integration with non-copyleft licenses, and Apache. These permissive licenses allow academic research and commercial use, making them fully compatible with the chosen Apache 2.0 license. Table˜6 lists all data currently included in NeuCo-Bench along with their origin. Google Earth Engine (GEE) [1] was utilized as the primary platform for downloading downstream task data as introduced in Section˜B.1. Future data and code contributions to NeuCo-Bench are required to be licensed under CC-BY 4.0 and Apache 2.0, respectively.
We note that the current NeuCo-Bench implementation lists CUDA packages covered by a proprietary NVIDIA license [^29] https://docs.nvidia.com/cuda/eula/index.html. However, we do neither bundle nor redistributes corresponding binaries. Users and contributors to NeuCo-Bench that share related docker containers need to explicitly attribute NVIDIA’s license. Fortunately, and as alluded in Table˜3 and Section˜A.2, NeuCo-Bench runs swiftly in a VM with commodity hardware specifications on CPU compute, only. Accordingly, the standalone NeuCo-Bench implementation introduced in Section˜B.2 can be started with (Bash) environment variable CUDA_VISIBLE_DEVICES=’’ to avoid usage of GPU resources.
Appendix C Baseline Methods
This appendix expands on the general evaluations introduced in the main paper by providing methodological details and extended results. Unlike the CVPR challenge setting, which separated development and evaluation splits, all results in the main paper and this appendix are computed on the full downstream datasets. Unless noted otherwise, we follow the evaluation protocol introduced in the main text, and use $E=20$ training epochs, $k=50$ train–test splits, and a learning rate of $10^{-3}$. We report raw $R^{2}$ values, clipping negative scores to $[0,1]$ only for visualization.
We begin by comparing the temporal aggregation methods in Section˜C.1 which motivate the use of post-encoding aggregation for all subsequent analyses. Thereafter, we report additional details of the baseline methods and results for the 1,024-dimensional embedding setup Section˜C.2. In Section˜C.3 we extend the ablations introduced in the main paper by providing per-task results for varying embedding dimensions and decoder choices beyond linear probing.
C.1 Temporal Aggregation Analysis